Emily invested $15,000 10 years ago at 8% compounded monthly. How much has she accumulated?
Answer #1

We use the formula:
A=P(1+r/12)^12n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=15000*(1+0.08/12)^(12*10)

=15000*2.21964023

=$33294.60(Approx).